OpenStudy (anonymous):

if 4^x = 7, what does 4^-2x equal?

OpenStudy (kc_kennylau):

Square the equation first :)

OpenStudy (kc_kennylau):

Then apply the reciprocal to the equation :)

OpenStudy (anonymous):

can you do it so i can see

OpenStudy (kc_kennylau):

What is the square of 4^x?

OpenStudy (kc_kennylau):

@ajprincess

OpenStudy (anonymous):

16^x^2?

OpenStudy (ajprincess):

4^(-2x)=4^(-2)*4^x 4^(-2)=1/4^2 4^2=4*4 the value of 4^x is given to u. so nw can u find the value of 4^(-2x)?

OpenStudy (anonymous):

nevermind. i understand the problem

OpenStudy (anonymous):

That x in 4^x looks like an infinity, lol\[4^x=7\]Square both sides.\[({4^x})^{2}=7^2\]When you have an exponent to another exponent, simply multiply them.\[4^{2x}=49\]Raise both sides to the -1 power. Again, we multiply exponents.\[4^{-2x}=49^{-1}\]The last step is to remember that raising something to a negative power means to take the reciprocal.\[4^{-2x}=\frac{1}{49}\]